Heart & Home Properties is developing a subdivision that includes 440 home lots. The 160 lots in the Canyon section are below a ridge and do not have views of the neighboring canyons and hills; the 280 lots in the Hilltop section offer unobstructed views. The expected selling price for each Canyon lot is $52,000 and for each Hilltop lot is $106,000. The developer acquired the land for $2,200,000 and spent another $2,300,000 on street and utilities improvements. Assign the joint land and improvement costs to the lots using the value basis of allocation and determine the average cost per lot. (Do not round your intermediate calculations.)
Answer #1
Average lot cost
Canyon section $              6,158
Hiltop section $           12,553
Workings:
Market Value Percentage of Market Value Cost to allocate Allocated cost Quantity of lots Average lot cost
Numerator Denominator % of Mkt Value
(a) (b) (c) (d) = (b) / (c) (e) (f) = (d) X (e) (g) (h) = (f) / (g)
Canyon section $     83,20,000 $      83,20,000 $ 3,80,00,000 21.89% $   45,00,000 $     9,85,263 160 $            6,158
Hiltop section $ 2,96,80,000 $ 2,96,80,000 $ 3,80,00,000 78.11% $   45,00,000 $   35,14,737 280 $         12,553
Totals $ 3,80,00,000 100.00% $   45,00,000